# Systems Security Foundations for Agentic Computing

### v1

@@ 5 Open Research Problems
We split the open problems into two categories. First we discuss three specific mechanisms that, if reliable and trustworthy, can solve a significant number of the security and privacy issues facing agentic deployments today. Second we discuss longer-term, fundamental topics that will allow for stronger security and privacy guarantees in future agentic deployments.
5.1 Mechanisms for Separating Instructions and Data
@@
One popular definition is that the agent/LLM should not follow any instructions that appear in the context window tagged as “data” [greshake2023notwhatyouvesignedupfor, wallace2024instructionhierarchytrainingllms, chen2024struqdefendingpromptinjection, chen2025secalign, wu2025instructionalsegmentembeddingimproving, chen2025defendingpromptinjectiondefensivetokens]. This can be restrictive in many tasks, as one of the key benefits of agents is to act on new information to complete an (often underspecified) goal. Learning to use an API and following web links based on information present in a web page are crucial to such adaptability, though they involve some form of following new instructions present in data. One option to introduce some flexibility into the instruction-following policy is to generalize the instruction-data separation and introduce “trust layers,” where system instructions are highest trust, user instructions are one level below, and instructions from tool data are even lower [wu2025instructionalsegmentembeddingimproving, wallace2024instructionhierarchytrainingllms, chen2025secalign]. If there is conflict of instructions, this priority is used to resolve. Unfortunately any nuanced policy requires the agent (or some additional tool) to correctly identify the new instructions inside the data.
These implementations do not offer security guarantees and, in this sense, are best effort. Attacks have been shown in the blackbox and white box setting [pandya2025iattentionbreakingfinetuning, wunderwuzzi2024breaking, nasr2025attackermovessecondstronger], as one can trick the model into following instructions despite the fine-tuning to learn separators. Furthermore, with multi-modal becoming standard, instructions can appear in images, video, audio, not just text. The past 10 years of adversarial machine learning has shown that “continuous domain” adversarial examples are easy to carry out [athalye2018obfuscatedgradientsfalsesense]. As a result, a probabilistic separation of instructions and data allows the attacker to employ an optimizer to easily find counter-examples to break the defense.
@@
For example, in an MCP tool poisoning attack, where the metadata in tool descriptions provides “canonical examples” of how a tool should be used [beurerkellner2025mcp, wang2025mindguard], an agent can be biased to make mistakes by generating tool calls using those canonical examples. Even if the agent does not follow an instruction, the arguments to tools its decides to use can still be poisoned by external sources. This follows from first principles: if the model is to “act” on data, then any of its subsequent outputs must necessarily reflect the influence of that data.
@@
5.2 Mechanisms for Access Control and Least Privilege
@@
The LLM/agent space does not follow any of these principles as, for example, an agent typically has uniform access to all tool calls, whether necessary for the current task or not. There are technical challenges to employing a least-privilege model. First, unlike traditional computer systems that had the notion of a program/application, for LLMs, there is no program. Rather, the LLM represents all possible programs at once because it can “compute” many different kinds of functions. Second, there is no “developer” to set access control policies, only a user who prompts the agent with a natural language task and who may lack any security expertise. Current recommendations from AI-agent providers highlight the need to proactively enable tools only as needed, as a form of manually enforced least privilege. For example OpenAI states in their documentation:
“Enable only the connectors needed for the current task.”
— from https://help.openai.com/en/articles/11752874-chatgpt-agent
A related approach that Google Gemini takes is to disable some of the tools available to the agent under certain conditions, but this is bypassable by a determined attacker that injects multiple layers of instructions [wunderwuzzi2024gemini].
There are current efforts at addressing the policy enforcement challenge for discrete tool-using agents where tools have proper semantic definitions [debenedetti2025defeatingpromptinjectionsdesign, progent2025privilege, li2026ace, syros2026saga]. These approaches provide a flexible way to express fine-grained access control, but there are AI agents (especially for computer-, browser-, phone-use tasks) that are not amenable to such system designs. Existing work also does not solve the policy specification problem. A complete access-control system also needs to provide tools for policy creation, maybe by inferring from the user’s prompt [10.1145/3713082.3730378] or by having system administrators in an enterprise write a mandatory access control policy. An interesting research opportunity is to use the reasoning capabilities of modern language models to create policy assistants that can have conversations with users to clarify gaps in natural language task specifications or to automatically decide when a user should be prompted with a permission screen.
@@
5.3 Mechanisms for Information-Flow Control
Access control as described above is a “gate” because it only decides a yes/no answer on whether the agent should have access to resources (one or more tools). There are many cases where an agent legitimately needs continued access to sensitive resources. For example, a coding agent might need access to API keys to perform operations like uploading a Docker container image to an endpoint. In such cases, the access-control policy will say that the agent has a legitimate reason to access the sensitive information. The least-privilege requirement is that the agent must use that information in a very specific way—in our example, to only upload the Docker image to the endpoint and nothing else. Thus, there is a need to ensure that the only allowed flow of information (the API key) is from the agent to the deployment endpoint. Information-flow control (IFC) is a well-researched primitive in the computer security literature but applying it to the LLM/agent space is challenging.
IFC works by labeling data, then tracking those labels, and enforcing label-based policies as the program operates on the data [tiwari2024ifc, taintdroid, denning-lattice]. This labeling and tracking can be done at many granularities (e.g., processor-instruction level, program-variable level, process level, filesystem-level, cross-computer level), with various guarantees. However, the common assumption is that it is possible to track labels as the corresponding data makes its way through the system and updating the labels accordingly, through “flow arithmetic.” [denning-lattice].
@@ 5.4 Long-Term: Security Guarantees from Probabilistic TCBs
are trying to abuse the system. Another particular complication, when using AI agents as part of the probabilistic TCB, is that agents may counterintuively try to evade the security policy due to their propensity to reward hacking [bondarenko2025demonstrating].
5.5 Long-Term: Security-Aware Model Architectures
The common approaches discussed up to now either place security mechanisms outside the agent or train the agent to perform the security task themselves. Another alternative may be to “surgically enhance” an already trained agent to enforce some security policy. If a circuit in the model is found to address security-relevant aspect of a task and the security policy blocks that aspect under some condition, it may be possible to enhance that circuit to discard its outputs when the condition holds [zou2024improvingalignmentrobustnesscircuit, hung2025attentiontrackerdetectingprompt]. Such a model would necessarily have an architecture that can be inspected (e.g., via mechanistic-interpretability means) and supplemented with security policy-specific constraints. While such an architecture may be more complex, it adds flexibility in allowing any security policy to be attached to the inference process, without having to train the model.
@@ 6 Conclusion
Information security and cryptography are classic fields with several well investigated principles and techniques. Provable defenses generally work by having an invariant called the security assumption (e.g., in case of system security a trusted computed base (TCB) that the attacker cannot tamper with, and in case of cryptography, the hardness of a mathematical problem, such as factoring or discrete log). In the context of AI-agents it is unclear what the security assumption could be. Without precise security assumptions, it is very challenging to build provable defenses. Another important principle in the systems security approach is to have programmable and deterministic guardrails at different layers of abstraction (separate instructions/data at the lowest level, perform least privilege access control at the system call/tool level and monitor/enforce information flows at the program/agent level). Applying the above-mentioned principle in the context of AI agents is challenging because the layers of abstractions for agentic systems are not clear. Traditional security principles and techniques are applicable and can thwart some attacks on agentic systems, but some attacks require us to solve new problems. However, new problems that emerge while applying traditional security principles and techniques to agentic systems provides an exciting agenda for the research community.

### v2

@@ 5 Open Research Problems
We split the open problems into two categories. First we discuss three specific mechanisms that, if reliable and trustworthy, can solve a significant number of security and privacy issues facing agentic deployments today. Second we discuss three longer-term, fundamental topics that will allow stronger security and privacy guarantees in future agentic deployments.
5.1 Separating Instructions and Data
@@
One popular definition is that the agent/LLM should not follow any instructions that appear in the context window tagged as “data” [39, 96, 16, 18, 102, 17]. However, this definition might be too restrictive in various settings, as one of the key benefits of agents is their ability to act upon new information to complete an (often underspecified) goal. Learning to use an API and following web links based on information present in a web page are crucial to such adaptability, though they involve some form of following new instructions present in data. One option to introduce some flexibility into the instruction-following policy is to generalize the instruction-data separation and introduce “trust layers,” where system instructions have the highest level of trust, whereas user instructions are one level below, and instructions from tool data have the lowest trust level [102, 96, 18]. If there is conflict of instructions, this priority is used to resolve it. Unfortunately, any nuanced policy requires the agent (or some additional tool) to explicitly and correctly identify the new instructions inside the data.
These implementations do not offer security guarantees and, in this sense, are best effort heuristics. Attacks have been shown in the black-box and white-box setting [65, 70, 60], as one can trick the model into following instructions despite the fine-tuning to learn separators. Furthermore, with multi-modal becoming standard, instructions may appear not only in text, but also, in images, video, and audio. The past 10 years of adversarial machine learning has shown that “continuous domain” adversarial examples are easy to carry out [7]. As a result, a probabilistic separation of instructions and data allows the attacker to employ an optimizer to easily find counter-examples to break the defense.
@@
For example, in an MCP tool poisoning attack, where the metadata in the tool descriptions provides “canonical examples” of how a tool should be used [10, 98], an agent can be biased to make mistakes by generating tool calls using those canonical examples. Even if the agent does not follow an instruction, the arguments to the tools that it decides to use can still be poisoned by external sources. This follows from first principles: if the model is to “act” on data, then any of its subsequent outputs must necessarily reflect the influence of that data.
@@
5.2 Access Control and Least Privilege
@@
The LLM/agent space does not follow any of these principles as, for example, an agent typically has uniform access to all tool calls, whether necessary for the current task or not. There are technical challenges to employing a least-privilege model. First, unlike traditional computer systems that had the notion of a program/application, for LLMs, there is no program. Rather, the LLM represents all possible programs at once because it can “compute” many different kinds of functions. Second, there is no “developer” to set access control policies, only a user who prompts the agent with a natural language task and who may lack any security expertise. Current recommendations from AI-agent providers highlight the need to proactively enable tools only as needed, as a form of manually enforced least privilege. For example OpenAI states in their documentation [63]: “Enable only the connectors needed for the current task.” A related approach that Google Gemini takes is to disable some of the tools available to the agent under certain conditions, but this is bypassable by a determined attacker that injects multiple layers of instructions [72].
There are current efforts at addressing the policy enforcement challenge for discrete tool-using agents where tools have proper semantic definitions [25, 84, 48, 90]. These approaches provide a flexible way to express fine-grained access control, but there are AI agents (especially for computer-, browser-, phone-use tasks) that are not amenable to such system designs. Existing work also does not solve the policy specification problem. A complete access-control system also needs to provide tools for policy creation, maybe by inferring from the user’s prompt [95] or by requiring system administrators in an enterprise write a mandatory access control policy. An interesting research opportunity is to use the reasoning capabilities of modern language models to create policy assistants that can have conversations with users to clarify gaps in natural language task specifications or to automatically decide when a user should be prompted with a permission screen.
@@
5.3 Information-Flow Control (IFC)
Access control as described above is a “gate” because it only decides a yes/no answer on whether the agent should have access to resources (one or more tools). There are many cases where an agent legitimately needs continued access to sensitive resources. For example, a coding agent might need access to API keys to perform operations like uploading a Docker container image to an endpoint. In such cases, the access-control policy will say that the agent has a legitimate reason to access the sensitive information. The least-privilege requirement is that the agent must use that information in a very specific way—in our example, to only upload the Docker image to the endpoint and nothing else. Thus, there is a need to ensure that the only allowed flow of information (the API key) is from the agent to the deployment endpoint. IFC is a well-researched primitive in the computer security literature but applying it to the LLM/agent space is challenging.
IFC works by labeling data, then tracking those labels, and enforcing label-based policies as the program operates on the data [92, 30, 27]. This labeling and tracking can be done at many granularities (e.g., processor-instruction level, program-variable level, process level, filesystem-level, cross-computer level), with various guarantees. However, the common assumption is that it is possible to track labels as the corresponding data makes its way through the system and updating the labels accordingly, through “flow arithmetic.” [27].
@@ 5.4 Long-Term: Security Guarantees from Probabilistic TCBs
are trying to abuse the system. Another particular complication, when using AI agents as part of the probabilistic TCB, is that agents may counterintuively try to evade the security policy due to their propensity to reward hacking [12].
5.5 Long-Term: Security-Aware ML Model Architectures
The common approaches discussed up to now either place security mechanisms outside the agent or train the agent to perform the security task themselves. Another alternative may be to “surgically enhance” an already trained agent to enforce some security policy. If a circuit in the model is found to address a security-relevant aspect of a task and the security policy blocks that aspect under some condition, it may be possible to enhance that circuit to discard its outputs when the condition holds [114, 42]. Such a model would necessarily have an architecture that can be inspected (e.g., via the means of mechanistic interpretability) and supplemented with security policy-specific constraints. While such an architecture may be more complex, it adds flexibility in allowing any security policy to be attached to the inference process, without having to train the model.
5.6 Long Term: Designing Correct Security Principles for LLMs
We have seen so far that Large Language Models (LLMs) by their very nature present an interesting dilemma. As we saw in Section 5.4, they are not completely deterministic computer programs, and they cannot be treated as such. They are like humans in many ways—they are probabilistic, they make mistakes, and often the mistakes are human-like because their training data is primarily generated by humans. Yet, there are also many aspects where they are not quite like humans—an LLM has access to vast information that no single human could ever aspire to. The possibility of solutions of a different nature thus arises: given a task and a situation, the LLM could reason or look through its vast body of knowledge to see if completing this task could result in a security failure.
All of this leads to an intriguing question: what are the right security principles while working with an LLM? What are some cases where we can treat it as a computer program (as in Section 5.4), where can we treat it as a human prone to mistakes, and where should we exploit its knowledge and reasoning ability to make better security decisions? Together with designing better probabilistic TCBs, this is yet another long-term question in building secure and effective agentic systems.
@@ 6 Conclusion
Information security and cryptography are classic fields with several well investigated principles and techniques. Provable defenses generally work by having an invariant called the security assumption. For example, in case of system security a TCB that the attacker cannot tamper with, and in case of cryptography, the hardness of a mathematical problem, such as factoring or discrete log. In the context of AI-agents it is unclear what the security assumption could be. Without precise security assumptions, it is very challenging to build provable defenses. Another important principle in the systems security approach is to have programmable and deterministic guardrails at different layers of abstraction (separate instructions/data at the lowest level, perform least privilege access control at the system call/tool level and monitor/enforce information flows at the program/agent level).
Applying the above-mentioned principle to AI agents is challenging because the abstraction layers in agentic systems are not yet well defined. Traditional security principles and techniques still apply and can thwart many attacks on agentic systems, but some attack classes expose genuinely new problems that require new solutions. Encouragingly, these emerging challenges create a compelling research agenda for the community.
Appendix A Example: Claude Code Data Exfiltration
Appendix B Security Principles and Mechanisms
Figure 3: Standard security architecture. Requests and responses cross a security boundary between an untrusted system and the trusted computing base (TCB). The TCB consults the security policy to decide whether it is permitted to answer a given request.
The standard security architecture consists of the trusted computing base (TCB), a security policy, the security boundary, and the untrusted system, as shown in Figure 3. The TCB consists of the functionality (code and data) whose integrity and confidentiality cannot be impacted by an attacker—in other words, the TCB defines the parts of the overall system that can be trusted to operate correctly under attack. The TCB typically contains the core functionality of the system (e.g., the kernel of the operating system) as well as a reference monitor that serves to approve or reject each request from the untrusted components to the TCB. The reference monitor is the part of the TCB that examines the security policy in order to make the approve/reject determination. The security policy is typically a declarative expression of the security goals—as an example, the security policy for stored files is given as a list of access control entries, each entry representing the operation(s) a given user is allowed to perform on a given file. The security boundary represents the interface through which the untrusted components interact with the TCB. In an OS, the security boundary between the OS kernel and the (untrusted) applications is the system-call interface.
Within this context we define and describe the security principles referred to in the case studies of Section 3.
Principle of Least Privilege.
This principle dictates that the Security Policy should only grant the Untrusted System the absolute minimum permissions it needs to function. For example, if the Untrusted System only needs to read a specific piece of data, the Security Policy should explicitly deny it permission to write, delete, or access any other data. The TCB is responsible for enforcing this minimal set of permissions, as specified by the Security Policy. For convenience most policies are written with the expectation of a default-deny fallback, where access to a resource not explicitly listed in the policy is automatically rejected.
TCB Tamper Resistance.
This principle states that the TCB itself must be protected from modification by any outside influence. The Security Boundary must be designed to prevent the Untrusted System from altering the code or logic of the TCB or the Security Policy. If the TCB could be tampered with, an attacker could disable the “Check for permission” and bypass all security controls.
Complete Mediation.
This principle dictates that every single request from the Untrusted System must be validated against the Security Policy. The system must not “remember” a previous authorization as users, trust levels, or other contextual conditions may have changed. Each time one of the requests crosses the Security Boundary, it must be intercepted by the TCB, which in turn must perform the “Check for permission to answer request” operation against the Security Policy. No request is allowed to bypass this check.
Secure Information Flow.
This principle governs the “Requests & Responses” channel, ensuring that sensitive information does not leak to untrusted areas. The Security Policy must define what kind of information is allowed to flow in which direction. For instance, even if a request is valid, the TCB must check the Security Policy to ensure that the resulting Response does not contain secret data that the Untrusted System is not authorized to see (this would constitute an unauthorized flow of data from high-trust components to low-trust components).
Human Weak Link.
This principle states that human operators can compromise this architecture in several ways, and thus the security mechanisms must be designed with this in mind. As a user, a human operating the Untrusted System could send malicious requests (e.g., SQL injection, prompt injection). As an administrator, a human could incorrectly configure the Security Policy, making it too permissive (violating Least Privilege). As a developer, a human could accidentally introduce a bug into the TCB that fails to perform the Check for permission operation correctly (violating Complete Mediation).